# Surrogate-Assisted Targeted Learning for Nested Bridge Functionals under Administrative Censoring

### 15 Mar 2026 v2

@@ Abstract
We develop a surrogate-assisted targeted minimum loss estimator (SA-TMLE) that resolves this obstruction via a two-stage nested fluctuation step, and establish three structural results: (i) under surrogate-mediated missing at random, the censoring mechanism contributes no separate tangent-space component to the efficient influence function; (ii) the standard DML one-step remainder contains $R_{SY}$ as an irreducible second-order term (Proposition 1); and (iii) the two-stage targeting step absorbs $R_{SY}$ into the efficient score without estimating $f_{S}$, yielding an asymptotically linear, doubly robust estimator under a single product-rate condition (Theorem 2). For clustered data, valid inference requires summation rather than averaging of individual influence contributions at the cluster level (Lemma 2); a leave-one-cluster-out jackknife applied to the full estimator restores near-nominal finite-sample coverage where the first-order sandwich interval undercovers.
Stepped-wedge cluster-randomized trials, in which calendar-driven rollout creates near-boundary observation probabilities for late-crossing clusters, provide the sharpest instance of this functional class and serve as the running example. Simulations confirm near-zero bias and jackknife coverage of 0.956–0.968 across all cluster sizes evaluated, against a backdrop where inverse-weighted competitors break down entirely at high censoring rates.
@@ 1 Introduction
A class of causal functionals in modern longitudinal studies shares the following semiparametric structure. The target parameter — an average treatment effect — is identified not by inverse-weighting incomplete observations, but by integrating an outcome regression over the conditional distribution of an intermediate variable. We call this a nested bridge functional: its efficient score has a cross-product structure that creates a specific failure mode for standard first-order estimators, and resolving that failure mode requires a targeting step that goes beyond ordinary debiased machine learning.
The observational setting that generates this structure is one where a primary endpoint requires substantial follow-up to realize, so that only units enrolled early enough have a fully matured observation by the time of analysis. A shorter-term surrogate is observed broadly across the study population regardless of enrollment timing. Rather than inverse-weighting the incomplete observations by $g_{\Delta}^{-1}$ — which becomes unstable whenever $g_{\Delta}$ concentrates near zero — one identifies the average treatment effect by integrating the complete-case outcome regression $\bar{Q}_{Y}(S,a,W,t)\equiv\mathbb{E}[Y\mid S,A=a,W,t,\Delta=1]$ over the treatment-specific conditional distribution of the surrogate:
@@
This representation requires only that $\bar{Q}_{Y}$ be estimable on the observed sub-cohort $\{\Delta=1\}$, not that $g_{\Delta}$ be bounded away from zero. The price is a new semiparametric complication: the functional (1) is nested in the sense that the outer integration is over the conditional surrogate law $f_{S}$, and the estimator must marginalize over $f_{S}$ without placing it in a denominator.
The core semiparametric obstruction.
The nested structure of (1) creates a specific failure mode for standard debiased machine-learning (DML) estimators. A natural one-step construction estimates $\bar{Q}_{\mathrm{int}}(a,W,t)\equiv\int\bar{Q}_{Y}(s,a,W,t)\,f_{S}(s\mid a,W,t)\,d\mu(s)$ by plugging in $\hat{\bar{Q}}_{Y}$ and $\hat{f}_{S}$ separately, leaving a second-order cross-product remainder
@@
Unlike the standard doubly-robust remainder, $R_{SY}$ has no complement in the efficient influence function: no nuisance parameter whose consistent estimation forces $R_{SY}=0$. DML cross-fitting eliminates first-order empirical-process terms but leaves $R_{SY}$ unchanged, since $R_{SY}$ is a $P_{0}$-expectation of a second-order product. Controlling $R_{SY}=o_{P}(J^{-1/2})$ therefore requires either a rate condition on $\hat{f}_{S}$ or an estimator that avoids placing $\hat{f}_{S}$ in the expansion. The SA-TMLE takes the second path: its two-stage nested fluctuation enforces $N^{-1}\sum D_{S,a}(O_{ijt})=0$, constraining $\hat{\bar{Q}}^{*}_{\mathrm{int}}$ to a propensity-weighted empirical mean of $\hat{\bar{Q}}^{*}_{Y}$ over observed surrogate values, absorbing $R_{SY}$ without ever estimating $f_{S}$.
Motivating instance.
Stepped-wedge cluster-randomized trials (SW-CRTs) provide a sharp instance of this structure. Calendar-determined treatment rollout means late-crossing clusters accumulate few primary-outcome observations before the administrative end date, creating a design-induced near-boundary regime $g_{\Delta}\to 0$ for those clusters. A short-term surrogate observed at crossover is available for all clusters, making the surrogate-bridge framework directly applicable. We use the SW-CRT as the running example throughout; the framework is not tied to this design.
Limitations of standard approaches.
IPCW estimators (Robins et al., 1994; Bang and Robins, 2005) place $g_{\Delta}^{-1}$ directly in the estimating equation, producing severe variance inflation as $g_{\Delta}\to 0$ and potentially failing to admit regular asymptotic approximations near the boundary. Parametric mixed models (Hussey and Hughes, 2007; Hemming and Taljaard, 2020; Hughes et al., 2020) restrict attention to the complete-case sub-cohort and require correct specification of the secular time trend and censoring mechanism (Kasza et al., 2019).
@@
This paper makes three contributions to semiparametric estimation theory, instantiated in the SW-CRT setting.
@@
Third, we construct the SA-TMLE and prove that its two-stage nested fluctuation absorbs $R_{SY}$ without estimating $f_{S}$, achieving $\sqrt{J}$-consistent asymptotic linearity under a single product-rate condition (Theorem 2). For finite-sample inference, a leave-one-cluster-out jackknife applied to the full estimator captures the fluctuation-step variance that the first-order sandwich misses, restoring near-nominal coverage across the cluster-count range evaluated.
Related statistical work.
Classical AIPW theory (Robins et al., 1994; Bickel et al., 1993) characterizes efficient estimation with missing outcomes but does not address the nested bridge structure or the $R_{SY}$ obstruction. The surrogate outcome literature (Gilbert et al., 2008; Frangakis and Rubin, 2002) addresses the complementary problem of rare (not delayed) primary outcomes; their identification strategies do not extend to the near-boundary censoring regime. Luedtke et al. (2017) construct a semiparametric surrogate-index estimator for i.i.d. data structurally related to ours, but their asymptotic linearity requires a uniform positivity lower bound on $g_{\Delta}$ that the present approach does not need. The DML framework (Chernozhukov et al., 2018) provides product-rate guarantees via sample splitting, but cross-fitting does not eliminate $R_{SY}$; this paper isolates that remainder (Proposition 1) and resolves it via a second targeting step. [^1] See Kennedy (2022) for a review of debiased ML. Cluster-level TMLE (Balzer et al., 2016, 2021) establishes clusterwise influence aggregation for group-randomized trials without the nested bridge structure studied here.
Section 2 defines the observed data structure and establishes non-parametric identification via the surrogate-bridge representation. Section 3 develops the semiparametric theory, including the EIC decomposition, the vanishing censoring-mechanism component (Lemma 1), the cluster-level summation rule (Lemma 2), and Proposition 1 characterizing $R_{SY}$. Section 4 constructs the SA-TMLE. Section 5 establishes asymptotic linearity, double robustness, and the Berry–Esseen coverage bound. Section 6 reports Monte Carlo experiments. Section 7 provides the design-calibrated EPT illustration.
@@ 2 Data Structure and Identification
$A_{jt}$ $=I(t\geq\tau_{j}),\quad\tau_{j}=f_{A}(U_{A,j}),$ (treatment; deterministic given $\tau_{j}$) (5)
@@ 3.1 EIC Decomposition
The individual-level EIC components, whose derivation is provided in Supplement, Section S2, are as follows. For treatment $a\in\{0,1\}$:
@@
where $\hat{\Psi}_{a}=N^{-1}\sum_{j,i}\bar{Q}^{\,*}_{\mathrm{int}}(a,W_{ijt},t_{ij})$.
@@
Collecting Proposition 3.1 for $a=1$ minus $a=0$:
@@ 5.2 Asymptotic Normality
The cluster-robust sandwich estimator (32) is consistent for $\sigma^{2}/J$, so the 95% confidence interval (33) achieves asymptotically exact coverage.
@@
Because $\bar{Q}^{\,*}_{\mathrm{int}}$ is thereby defined as a propensity-weighted empirical average of $\bar{Q}^{\,*}_{Y}$ over the observed surrogate distribution—rather than as an explicit integral against an estimated $\hat{f}_{S}$—the cross-product $(\hat{Q}_{Y}-Q^{0}_{Y})(\hat{f}_{S}-f^{0}_{S})$ that constitutes $R_{SY}$ never enters the expansion; $R_{SY}$ is absorbed into the efficient score without any requirement on the rate of $\hat{f}_{S}$. The formal calculation is in Supplement, Section S2.6.
@@
Asymptotic normality. With $R_{2}+R_{SY}=o_{P}(J^{-1/2})$ established, the Lindeberg–Feller CLT applied to the i.i.d. cluster scores $\{\mathrm{EIC}_{j}\}_{j=1}^{J}$ gives (37). Consistency of the sandwich estimator (32) for $\sigma^{2}/J$ follows by the law of large numbers applied at the cluster level. The complete argument is in Supplement, Section S2.4.∎∎
@@ 5.3 Non-asymptotic Coverage Bound
The leave-one-cluster-out jackknife for the full SA-TMLE estimator resolves this gap in practice (Section 6.3), yielding near-nominal coverage at $J\geq 30$. Formal jackknife variance consistency for targeted estimators with data-adaptive nuisance follows from the arguments of Zheng and van der Laan (2011); a complete non-asymptotic treatment is deferred to future work.
@@ 6.3 Results
Jackknife variance and coverage restoration. The leave-one-cluster-out jackknife, applied to the full two-stage estimator (re-running both fluctuation steps for each held-out cluster), captures the fluctuation-step variance that the sandwich misses. Table 1 shows that the mean jackknife variance tracks the empirical variance closely at all $J$. The resulting coverage (Table 2, Jackknife columns) is near-nominal at the practically important cluster sizes: 0.956 at $J=30$ and 0.960 at $J=50$. At very small $J$ (10 and 20) the jackknife is mildly conservative (0.970–0.980), as the leave-one-out procedure loses a larger fraction of training data per fold and slightly overestimates variability; this is a well-understood property of the jackknife in small samples and represents a safe failure mode. By $J\geq 30$ the conservatism resolves and coverage is stable near 0.956–0.968 through $J=100$.
The power tradeoff is modest: at $J=30$, jackknife power is 0.57 versus sandwich power of 0.73, reflecting wider intervals. By $J=50$, jackknife power reaches 0.786, and by $J=100$ it is 0.970 — essentially the same as the sandwich at those sizes. The undercoverage was a variance-calibration issue, not a failure of the point estimator; the cluster jackknife resolves it without altering the bias or RMSE.
@@
Table 3 reports performance at $J=30$ under the three misspecification regimes of Section 6.2. Under scenario (ii) (outcome models correct, $g_{\Delta}$ misspecified), CV-TMLE achieves near-zero bias (<0.001<0.001) and coverage of 0.921, providing supportive finite-sample evidence for the asymptotic double-robustness result in the regime where the outcome regressions are correctly specified: the correctly specified outcome models drive the product-rate condition (35) toward $o_{P}(J^{-1/2})$, consistent with Remark 3.
Under scenario (i) (outcome models misspecified, $g_{\Delta}$ correct), CV-TMLE exhibits bias of 0.100 with coverage 0.426. This illustrates an important finite-sample limitation of the asymptotic double-robustness guarantee: when the misspecified outcome regression leaves a large structured error, correct specification of $g_{\Delta}$ alone may not be sufficient to deliver good finite-sample performance at $J=30$. In this setting, the targeting step appears to reduce but not eliminate the residual bias induced by the misspecified outcome model.
@@
Table 3: Block II simulation results: double robustness under nuisance misspecification ($J=30$, $\Psi^{*}=0.12$, 1,000 replicates). Scenario labels: (i) outcome models only misspecified; (ii) censoring propensity only misspecified; (iii) both misspecified.
@@
Additional simulations varying $\sigma_{b}^{2}\in\{0.006,0.034,0.071,0.160\}$ (ICC $\approx$ 0.01–0.20) confirm that the jackknife restores near-nominal coverage (0.946–0.968) across the full empirically observed ICC range in implementation-science SW-CRTs, extending the Block I finding to the ICC heterogeneity setting; full results are in Supplement, Section S4.

### 31 Mar 2026 v4

@@ Abstract
We develop a surrogate-assisted targeted minimum loss estimator (SA-TMLE) that circumvents this obstruction via a two-stage nested fluctuation step, and establish three structural results: (i) under surrogate-mediated missing at random, the censoring mechanism contributes no separate tangent-space component to the efficient influence function; (ii) the standard DML one-step remainder contains $R_{SY}$ as an irreducible second-order term (Proposition 1); and (iii) the two-stage targeting step absorbs $R_{SY}$ into the efficient score without estimating $f_{S}$, yielding an asymptotically linear, doubly robust estimator under a single product-rate condition (Theorem 2). For clustered data, valid inference requires summation rather than averaging of individual influence contributions at the cluster level (Lemma 2); a leave-one-cluster-out jackknife applied to the full estimator is proved variance-consistent under stated regularity conditions (Theorem 3) and yields near-nominal finite-sample coverage in the simulation regimes studied, where the first-order sandwich interval undercovers.
Stepped-wedge cluster-randomized trials, in which calendar-driven rollout creates near-boundary observation probabilities for late-crossing clusters, provide the sharpest instance of this functional class and serve as the running example. Simulations confirm near-zero bias and jackknife coverage of 0.956–0.968 in the moderate-to-large cluster regimes studied, against a backdrop where inverse-weighted competitors break down entirely at high censoring rates.
@@ 1 Introduction
Consider an average treatment effect functional in which the primary outcome is missing for a non-negligible fraction of units, while an earlier intermediate variable — a surrogate — is broadly observed. Standard semiparametric estimators handle missingness by inverse-weighting the complete cases with $g_{\Delta}^{-1}$, the reciprocal observation probability. When $g_{\Delta}$ concentrates near zero, this approach produces severe variance inflation and may fail to admit a regular asymptotic linear representation. An alternative identification strategy avoids $g_{\Delta}^{-1}$ entirely: integrate the complete-case outcome regression $\bar{Q}_{Y}(S,a,W,t)\equiv\mathbb{E}[Y\mid S,A=a,W,t,\Delta=1]$ over the treatment-specific conditional distribution of the surrogate,
@@
We call (1) a nested bridge functional: the outer integration is over the conditional surrogate law $f_{S}$ rather than a propensity denominator, and the estimator must marginalize over $f_{S}$ without placing it in a denominator. This functional class arises generically whenever a delayed primary outcome is administratively censored and a short-term surrogate satisfies a conditional independence condition (Assumption 3). Instances include randomized trials with staggered enrollment and a fixed analysis date, registry studies with incomplete long-term follow-up, and — most sharply — stepped-wedge cluster-randomized trials (SW-CRTs), in which calendar-driven treatment rollout creates a design-induced near-boundary regime $g_{\Delta}\to 0$ for late-crossing clusters.
The nested structure of (1) creates a distinctive semiparametric obstruction. A natural one-step debiased machine-learning (DML) construction estimates the integrated regression $\bar{Q}_{\mathrm{int}}(a,W,t)\equiv\int\bar{Q}_{Y}(s,a,W,t)\,f_{S}(s\mid a,W,t)\,d\mu(s)$ by plugging in $\hat{\bar{Q}}_{Y}$ and $\hat{f}_{S}$ separately, leaving a second-order cross-product remainder
@@
Unlike the standard doubly-robust remainder, $R_{SY}$ has no complement in the efficient influence function: no nuisance parameter whose consistent estimation forces $R_{SY}=0$. DML cross-fitting eliminates first-order empirical-process terms but leaves $R_{SY}$ unchanged, since $R_{SY}$ is a $P_{0}$-expectation of a second-order product. Controlling $R_{SY}=o_{P}(J^{-1/2})$ therefore requires either a rate condition on $\hat{f}_{S}$ or an estimator that avoids placing $\hat{f}_{S}$ in the expansion altogether.
We develop a surrogate-assisted targeted minimum loss estimator (SA-TMLE) that takes the second path: its two-stage nested fluctuation enforces $N^{-1}\sum D_{S,a}(O_{ijt})=0$, constraining $\hat{\bar{Q}}^{*}_{\mathrm{int}}$ to a propensity-weighted empirical mean of $\hat{\bar{Q}}^{*}_{Y}$ over observed surrogate values, absorbing $R_{SY}$ without ever estimating $f_{S}$.
The stepped-wedge trial as sharpest instance.
SW-CRTs provide the sharpest concrete realization of the nested bridge class. Calendar-determined treatment rollout means late-crossing clusters accumulate few primary-outcome observations before the administrative end date, creating $g_{\Delta}\to 0$ by design for those clusters. A short-term surrogate observed at crossover is available for all clusters, making the surrogate-bridge functional directly applicable. We use the SW-CRT as the running example throughout to make identification conditions, positivity boundaries, and clustered-data complications concrete; the theory applies to any setting in which identification takes the form (1).
Limitations of existing approaches.
IPCW estimators (Robins et al., 1994; Bang and Robins, 2005) place $g_{\Delta}^{-1}$ directly in the estimating equation, producing severe variance inflation as $g_{\Delta}\to 0$ and potentially failing to admit regular asymptotic approximations near the boundary. Parametric mixed models (Hussey and Hughes, 2007; Hemming and Taljaard, 2020; Hughes et al., 2020) restrict attention to the complete-case sub-cohort and require correct specification of the secular time trend and censoring mechanism (Kasza et al., 2019). The DML framework (Chernozhukov et al., 2018) provides product-rate guarantees via sample splitting, but cross-fitting does not eliminate $R_{SY}$; this paper isolates that remainder (Proposition 2) and resolves it via a second targeting step. [^1] See Kennedy (2022) for a review of debiased ML.
@@
The main contribution of this paper is a targeted inference strategy for nested bridge functionals under surrogate-mediated administrative censoring that avoids inverse observation weights and remains reliable under clustered sampling. Three supporting results clarify why this construction is needed and how it behaves.
@@
Third, we construct the SA-TMLE and prove that its two-stage nested fluctuation absorbs $R_{SY}$ without estimating $f_{S}$, achieving $\sqrt{J}$-consistent asymptotic linearity under a single product-rate condition (Theorem 2). For finite-sample inference, a leave-one-cluster-out jackknife applied to the full estimator is shown to be variance-consistent under mild regularity conditions (Theorem 3), and yields near-nominal coverage across the cluster-count range evaluated in simulation.
Related work.
Classical AIPW theory (Robins et al., 1994; Bickel et al., 1993) characterizes efficient estimation with missing outcomes but does not address the nested bridge structure or the $R_{SY}$ obstruction. The surrogate outcome literature (Gilbert et al., 2008; Frangakis and Rubin, 2002) addresses the complementary problem of rare (not delayed) primary outcomes; their identification strategies do not extend to the near-boundary censoring regime. Athey et al. (2025) introduce the surrogate index — the conditional expectation of the primary outcome given short-term proxies — and establish identification under a statistical surrogacy condition; their semiparametric estimator requires a uniform positivity lower bound on $g_{\Delta}$ that the present approach does not need. Kallus and Mao (2025) derive efficiency bounds for ATE estimation when surrogate observations supplement limited primary-outcome data without imposing strong surrogacy; their doubly robust estimator occupies the same functional class as the nested bridge but does not address the $R_{SY}$ obstruction or the near-boundary censoring regime. Cluster-level TMLE (Balzer et al., 2016, 2021) establishes clusterwise influence aggregation for group-randomized trials without the nested bridge structure studied here.
Section 2 defines the observed data structure and establishes non-parametric identification via the surrogate-bridge representation. Section 3 develops the semiparametric theory, including the EIC decomposition, the vanishing censoring-mechanism component (Lemma 1), the cluster-level summation rule (Lemma 2), and Proposition 2 characterizing $R_{SY}$. Section 4 constructs the SA-TMLE. Section 5 establishes asymptotic linearity, double robustness, the Berry–Esseen coverage bound, and jackknife variance consistency. Section 6 reports Monte Carlo experiments. Section 7 provides the design-calibrated EPT illustration.
@@ 2 Data Structure and Identification
$A_{jt}$ $=I(t\geq\tau_{j}),\qquad\tau_{j}=f_{A}(U_{A,j}),$ (treatment; deterministic given $\tau_{j}$) (5)
@@ 3.1 EIC Decomposition
The individual-level EIC components are derived in Supplement, Section S2, via pathwise differentiation on each sub-tangent space; the result is stated here as a proposition for later reference.
Proposition 1 (Individual-Level EIC Components).
Under Assumptions 15, for treatment $a\in\{0,1\}$:
@@
where $\hat{\Psi}_{a}=N^{-1}\sum_{j,i}\bar{Q}^{\,*}_{\mathrm{int}}(a,W_{ijt},t_{ij})$. The total individual EIC is:
@@
Collecting Proposition 1 for $a=1$ minus $a=0$:
@@ 5.2 Asymptotic Normality
The cluster-robust sandwich estimator (32) is consistent for the first-order asymptotic variance $\sigma^{2}/J$. When the two-stage fluctuation parameters $(\varepsilon_{Y},\varepsilon_{S})$ have non-degenerate sampling variability, the total finite-sample variance exceeds $\sigma^{2}/J$ by an additive $O(J^{-1})$ term; the jackknife variance estimator of Theorem 3 below captures this additional component and delivers asymptotically exact coverage.
@@
Because $\bar{Q}^{\,*}_{\mathrm{int}}$ is thereby defined as a propensity-weighted empirical average of $\bar{Q}^{\,*}_{Y}$ over the observed surrogate distribution—rather than as an explicit integral against an estimated $\hat{f}_{S}$—the cross-product $(\hat{Q}_{Y}-Q^{0}_{Y})(\hat{f}_{S}-f^{0}_{S})$ that constitutes $R_{SY}$ never enters the expansion; $R_{SY}$ is absorbed into the efficient score without any requirement on the rate of $\hat{f}_{S}$. The formal calculation is in Supplement, Section S2.6. In practice, the score equation (39) is solved by a one-parameter quasibinomial GLM with logistic link, which achieves numerical convergence to machine precision (typically $10^{-8}$); the residual score contributes a negligible $O(10^{-8})$ term to $R_{SY}$.
@@
Asymptotic normality. With $R_{2}+R_{SY}=o_{P}(J^{-1/2})$ established, the Lindeberg–Feller CLT applied to the i.i.d. cluster scores $\{\mathrm{EIC}_{j}\}_{j=1}^{J}$ gives (37). Consistency of the sandwich estimator (32) for the first-order variance $\sigma^{2}/J$ follows by the law of large numbers applied at the cluster level; the gap between $\sigma^{2}/J$ and the total finite-sample variance is characterised in Theorem 3. The complete argument is in Supplement, Section S2.4.∎∎
@@ 5.3 Non-asymptotic Coverage Bound
The leave-one-cluster-out jackknife for the full SA-TMLE estimator resolves this gap in practice (Section 6.3), yielding near-nominal coverage at $J\geq 30$. The following theorem, proved in Supplement, Section S2.7, establishes that the jackknife is consistent for the total finite-sample variance of the SA-TMLE — including the fluctuation-step contribution that the sandwich misses — rather than merely the first-order asymptotic variance.
Theorem 3 (Jackknife Variance Consistency for the SA-TMLE).
Suppose Assumptions 15 and conditions (C1)–(C3) of Theorem 2 hold. Suppose additionally:
  1. (C5)
    (Jackknife stability) The nuisance estimators $\hat{\eta}^{(-j)}$ trained on J1J-1 clusters satisfy $\sup_{j}\|\hat{\eta}^{(-j)}-\hat{\eta}\|_{P_{0}}=O_{P}(J^{-1})$, and the fluctuation-step map $\eta\mapsto\varepsilon(\eta)$ is Lipschitz in a neighbourhood of $\eta_{0}$ with respect to the $L^{2}(P_{0})$ norm.
Define the leave-one-cluster-out jackknife variance estimator
$\hat{\sigma}^{2}_{\mathrm{jack}}\;=\;\frac{J-1}{J}\sum_{j=1}^{J}\bigl(\hat{\Psi}_{\mathrm{TMLE}}^{(-j)}-\bar{\Psi}^{(\cdot)}\bigr)^{2},$ (40)
where $\hat{\Psi}_{\mathrm{TMLE}}^{(-j)}$ is the full SA-TMLE (including both fluctuation steps) refit on the J1J-1 clusters excluding cluster $j$, and $\bar{\Psi}^{(\cdot)}=J^{-1}\sum_{j}\hat{\Psi}_{\mathrm{TMLE}}^{(-j)}$.
Write $\sigma^{2}_{\infty}=\mathrm{Var}(\mathrm{EIC}_{j})$ for the first-order asymptotic variance and $\sigma^{2}_{J}=J\cdot\mathrm{Var}(\hat{\Psi}_{\mathrm{TMLE}})$ for the exact (scaled) finite-sample variance. The two-stage targeting construction induces a decomposition
$\sigma^{2}_{J}\;=\;\sigma^{2}_{\infty}\;+\;V_{\varepsilon}\;+\;o(1),$ (41)
where $V_{\varepsilon}\geq 0$ is the variance contribution from the fluctuation-step parameters $(\varepsilon_{Y},\varepsilon_{S})$, which are re-estimated on each leave-one-out sample. Then:
  1. (a)
    The sandwich variance estimator (32) satisfies $\widehat{\mathrm{Var}}_{\mathrm{sand}}\to_{P}\sigma^{2}_{\infty}/J$ and is therefore consistent for the first-order asymptotic variance only.
  2. (b)
    The jackknife variance estimator satisfies
    $\hat{\sigma}^{2}_{\mathrm{jack}}\;=\;\frac{\sigma^{2}_{J}}{J}+o_{P}(J^{-1})\;=\;\frac{\sigma^{2}_{\infty}+V_{\varepsilon}}{J}+o_{P}(J^{-1}),$ (42)
    and is thus consistent for the total finite-sample variance.
When $V_{\varepsilon}>0$ — which holds whenever the fluctuation-step parameters have non-degenerate sampling variability, as is generic for the two-stage SA-TMLE — the sandwich strictly underestimates the true variance by the ratio $\sigma^{2}_{J}/\sigma^{2}_{\infty}>1$, and this ratio does not vanish as $J\to\infty$. The jackknife confidence interval $\hat{\Psi}_{\mathrm{TMLE}}\pm t_{J-1,\alpha/2}\,\hat{\sigma}_{\mathrm{jack}}$ achieves asymptotically exact $1-\alpha$ coverage.
Remark 4 (Reconciliation with Table 1).
The simulation ratio $\mathrm{Var}_{\mathrm{emp}}/\mathrm{Sand.}\approx 1.6$ across all $J$ is explained by (41): $V_{\varepsilon}/\sigma^{2}_{\infty}\approx 0.6$ at the simulation DGP, so $\sigma^{2}_{J}/\sigma^{2}_{\infty}\approx 1.6$. Because both $\sigma^{2}_{\infty}$ and $V_{\varepsilon}$ are $O(1)$, the ratio is asymptotically constant — precisely the pattern observed.
@@ 6.3 Results
Jackknife variance and coverage restoration. The leave-one-cluster-out jackknife, applied to the full two-stage estimator (re-running both fluctuation steps for each held-out cluster), captures the fluctuation-step variance that the sandwich misses. Table 1 shows that the mean jackknife variance tracks the empirical variance closely at all $J$. The resulting coverage (Table 2, Jackknife columns) is near-nominal at the practically important cluster sizes: 0.956 at $J=30$ and 0.960 at $J=50$. At very small $J$ (10 and 20) the jackknife is mildly conservative (0.970–0.980), as the leave-one-out procedure loses a larger fraction of training data per fold and slightly overestimates variability; this pattern is consistent with the finite-sample conservatism often seen for leave-one-out procedures when $J$ is small and represents a safe failure mode. By $J\geq 30$ the conservatism attenuates and coverage remains stable near 0.956–0.968 through $J=100$.
The power tradeoff is modest: at $J=30$, jackknife power is 0.57 versus sandwich power of 0.73, reflecting wider intervals. By $J=50$, jackknife power reaches 0.786, and by $J=100$ it is 0.970 — essentially the same as the sandwich at those sizes. The undercoverage was a variance-calibration issue, not a failure of the point estimator; the cluster jackknife substantially corrects this variance-calibration gap without materially altering the bias or RMSE.
@@
Table 3 reports performance at $J=30$ under the three misspecification regimes of Section 6.2. Under scenario (ii) (outcome models correct, $g_{\Delta}$ misspecified), CV-TMLE achieves near-zero bias (<0.001<0.001) and sandwich coverage of 0.921, providing supportive finite-sample evidence for the asymptotic double-robustness result in the regime where the outcome regressions are correctly specified: the correctly specified outcome models drive the product-rate condition (35) toward $o_{P}(J^{-1/2})$, consistent with Remark 3. The jackknife interval achieves 0.944 in this scenario, near-nominal and substantially above the sandwich, confirming that the variance-calibration gap is resolved by the full two-stage leave-one-out refitting even under propensity misspecification.
Under scenario (i) (outcome models misspecified, $g_{\Delta}$ correct), CV-TMLE exhibits bias of 0.100 with coverage 0.426. This quantifies the finite-sample cost of relying on the propensity arm of double robustness alone: the misspecified outcome model contributes $\|\hat{\bar{Q}}_{Y}-\bar{Q}^{0}_{Y}\|_{P_{0}}=\Theta(1)$ (a constant-order structured error), and although the correctly specified $g_{\Delta}$ achieves $\|\hat{g}_{\Delta}-g^{0}_{\Delta}\|_{P_{0}}=O_{P}(J^{-1/2})$, their product is $\Theta(J^{-1/2})$ — at the boundary of condition (C3), not strictly $o_{P}(J^{-1/2})$. At $J=30$, the resulting $R_{2}$ is $O_{P}(J^{-1/2})\approx 0.18$, which is commensurate with the observed bias of $0.100$. Asymptotic double robustness guarantees $R_{2}\to 0$, but the convergence is too slow for the product to be negligible at this $J$.
@@
Table 3: Block II simulation results: double robustness under nuisance misspecification ($J=30$, $\Psi^{*}=0.12$, 1,000 replicates). Scenario labels: (i) outcome models only misspecified; (ii) censoring propensity only misspecified; (iii) both misspecified. Coverage uses $t_{J-1}$ critical values with the cluster-robust sandwich variance estimator. Jackknife coverage for CV-TMLE under scenario (ii) is 0.944, consistent with near-nominal performance when outcome models are correctly specified.
@@
Additional simulations varying $\sigma_{b}^{2}\in\{0.006,0.034,0.071,0.160\}$ (ICC $\approx$ 0.01–0.20) show that the jackknife yields near-nominal coverage (0.946–0.968) across the empirically observed ICC range in implementation-science SW-CRTs, extending the Block I finding to the ICC heterogeneity setting; full results are in Supplement, Section S4.